OpenStudy (anonymous):

calculate d/dx f(x)= (x^2+1)^2-(x^2+1)

OpenStudy (anonymous):

2(2x)-2x i think

OpenStudy (anonymous):

make sure you multiply that first thing out

OpenStudy (anonymous):

2 x (2 x^2+1) = f'(x)

OpenStudy (anonymous):

x^4 +2x^2 + 1 - (x^2 +1) -----> d/dx = 4x^3 + 4x - 2x

OpenStudy (anonymous):

= 4x^3 +2x

OpenStudy (anonymous):

or x(4x^2 + 2) if you wanna be that guy

OpenStudy (anonymous):

f'(x) = 2*(x^2 +1)*2x - 2x = 4x(x^2+1) - 2x = 4x^3 + 4x - 2x = 4x^3 + 2x

OpenStudy (anonymous):

the chain rule is used for first part of function